Translate to an equation.
f=folder's at $0.30 apiece
and n=notebooks at $1.50 apiece
The total of folders and notebooks is $12
Graph the equation, and use the graph to determine three different combinations of folders and notebooks that total $12

Answer by fcabanski(1391)

12 = .3f + 1.5n


It's a linear equation, so find two points. Try when n=0 and f=0.


If n=0 (0 notebooks) then it's 12 = .3f * 0
12/.3 = f = 40


If f=0 (0 folders) then it's 12 = 0 + 1.5n
12/1.5 = n = 8.


Those points show two combinations (n,f): (8,0) and (0,40).


If you graph it and look at the line, you'll find a third whole number combination (4,20).

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com